120% of 500 miles is how many miles?
Found 2 solutions by stanbon, ale6chu:
Answer by stanbon(75887)

120% of 500 miles is how many miles?
---
120% = 120/100 = 1.2
---
Ans: 1.2*500 = 600
=====================
Cheers,
Stan H.
=====================

Answer by ale6chu(1)